# Log Parsing with Self-Generated In-Context Learning and Self-Correction

### v1

@@ 3.2. Tree-based Log Parser
In practical applications, the number of log templates is significantly smaller than the number of log messages (Jiang et al., 2023; Wang et al., 2022). For instance, the Loghub-2.0 dataset (Jiang et al., 2023) contains over 50 million logs but fewer than 3,500 templates. This disparity makes it inefficient to parse each log message individually using LLMs. To improve efficiency, AdaParser employs a tree-based parser, leveraging a trie-based parsing tree to store templates generated by the LLM and quickly match templates for new log messages. Each intermediate node in the tree represents a token from the log templates, with the wildcard “¡*¿” matching sequences of tokens of any length. The leaf node in the tree represents a complete log template, formed by concatenating all tokens along the path from the root to the leaf.
@@
When a new log message arrives, AdaParser tokenizes it and matches the tokens against the parsing tree. If the matching process reaches a leaf node, it indicates that the template stored at the leaf node exactly matches the log message, and AdaParser directly returns the template without further LLM querying. If the matching process terminates at an intermediate node, the templates within the subtrees of the intermediate node form a list of relevant templates, denoted as $[T_{1},T_{2},...,T_{n}]$, which share the same prefixes as the input log message. AdaParser then utilizes the SG-ICL-based Parser (Sec. 3.3) to parse the log message and generates a corrected template $T_{c}$ through the Template Corrector (Sec. 3.4). Subsequently, the parsing tree is updated based on the similarity between $T_{c}$ and the relevant templates. Specifically, we calculate the similarity between $T_{c}$ and each relevant template $T_{i}$ in $[T_{1},T_{2},...,T_{n}]$. We split them into lists of tokens, denoted as $L_{c}$ and $L_{i}$. The similarity between $L_{c}$ and $L_{i}$ is defined as follows:
(1) $sim(T_{c},T_{i})=\frac{2\times len(LCS(L_{c},L_{i}))}{len(L_{c})+len(L_{i})}$
@@
where LCS is the longest common subsequence of the two templates. If the similarity exceeds a predefined threshold (e.g., 0.8 in our implementation), it indicates that $T_{c}$ and $T_{i}$ are highly similar and likely derived from the same ground-truth template. In this case, AdaParser merges $T_{c}$ and $T_{i}$ by replacing the different tokens in the path of $T_{i}$ with the wildcard “¡*¿”. If the similarity is below the threshold, it means that $T_{c}$ has a low correlation with $T_{i}$. However, templates with low similarity can still originate from the same ground-truth template. For instance, in the Linux Dataset from Loghub-2.0, the similarity between the log message “apmd shutdown succeeded” and the relevant messages “ntpd shutdown succeeded”, “crond shutdown succeeded”, and “xinetd shutdown succeeded” is 0.67, which is below the threshold. Nonetheless, they all share the same template “¡*¿ shutdown succeeded”. To address this issue, we further calculate the number of different tokens in the same position among these templates with the same similarity. If this number exceeds a predefined threshold (e.g., 5 in our implementation), AdaParser merges $T_{c}$ and $T_{i}$. Otherwise, AdaParser inserts $T_{c}$ into the parsing tree.
@@
Figure 4. The workflow of SG-ICL-based parser. After step $t$, the message with the template generated by the LLM is added to the candidate set and then used as a demonstration at step $t+1$.
3.3. SG-ICL-based Parser
@@
The ICL paradigm initializes a candidate set by sampling a small set of log messages and corresponding templates from the historical log data. LLMs can leverage system-specific features within demonstrations from the candidate set to facilitate more accurate log parsing. We utilize the hierarchical sampling algorithm (JIANG et al., 2023) to extract a set of log messages from historical log data as the candidate set. However, as discussed in Sec. 2.3, sufficient log data is not always available. Therefore, AdaParser supports initializing an empty candidate set, which is incrementally populated by SG-ICL.
@@
Figure 4 illustrates the workflow of SG-ICL-based parser. At step 00, if the candidate set is empty, we utilize zero-shot prompting to generate templates. At step $t$, when a new log message arrives, we first select similar demonstrations from the candidate set. Several studies have shown that demonstrations similar to the query can help LLMs better understand the desired behavior (Nashid et al., 2023; Gao et al., 2023; Geng et al., 2024). Hence, we use a kNN-based demonstration selection algorithm that is computationally efficient in practice. Specifically, for a queried log message $q$ and all candidate log messages $c_{i}$, we tokenized $q$ and $c_{i}$ into lists of tokens using delimiters such as punctuations and whitespace. Then, we calculate the LCS similarity $sim(q,c_{i})$ as shown in Eq. 1 and select the top-k most similar logs from the candidate set as demonstrations. These logs, having similar tokens and special formats to the queried log, assist the LLM in understanding the semantics of the queried log.
@@
These demonstrations are assembled into the parsing prompt. Recent work (Zhao et al., 2021; Gao et al., 2023) has pinpointed that LLMs are more prone to be influenced by the demonstrations that are closer to the query. Hence, we arrange the demonstrations in ascending order of similarity to the queried log. This is based on the intuition that demonstrations with higher similarity may contain more information related to the queried log. AdaParser then utilizes this parsing prompt to query the LLM and obtain the generated templates. Subsequently, a new example - the current log message paired with the generated template - is added to the candidate set, facilitating continuous adaptation to evolving log data. At step $t+1$, the new example from the previous step $t$ becomes part of the candidate set and can be selected as a demonstration to help parse upcoming log messages.
3.4. Template Corrector
Despite the considerable empirical success of LLMs, these models are not devoid of flaws. They occasionally exhibit undesired and inconsistent behaviors, such as generating seemingly convincing but inaccurate “hallucinations” and failing to trustfully adhere to rules and constraints (Pan et al., 2024). In particular, when applied to log parsing, LLMs are vulnerable to two types of parsing errors, as detailed in Sec. 2.3. These errors can undermine parsing accuracy and pose obstacles for downstream tasks. A prevailing strategy to correct these undesired behaviors in LLMs involves learning from feedback, analogous to the typical human learning process where individuals refine their behaviors through cycles of trial, error, and correction. When making mistakes, humans often acquire feedback either from others or through self-reflection (Boyd and Fales, 1983; Metcalfe, 2017). Such feedback provides valuable insights that enable humans to correct mistakes and modify their behavior accordingly.
@@ 3.4. Template Corrector
Figure 5. Two examples of self-correction. Template matching correction and variable abstracting correction are used to correct the two types of parsing errors. The corrected parts are marked in red.
3.4.1. Template Matching Correction
@@
A correct template must accurately match the corresponding log message, otherwise, the template suffers from the error of Plausible template. To fix this error, we propose a template matching correction strategy. Specifically, given a log message and a template generated by the LLM, we first verify whether the generated template exactly matches the input log message. This verification involves transforming the generated template into a regular expression (e.g., replace “¡*¿” as “.*”) to check if it matches the input log message exactly. If the regular expression fails to match the input log message, we employ a designed correcting prompt to rectify the template. As illustrated in Figure 5a, the LLM initially parsed the token “0x300sent” incorrectly, treating it as a variable “0x300” and a constant “sent”. This misinterpretation led to a mismatch between the input log message and the generated log template. After querying the LLM to revise the log template, the LLM recognized that “0x300sent” should be treated as a variable and abstracted it with the wildcard “¡*¿”, thereby successfully correcting the error in the initially generated template.
3.4.2. Variable Abstracting Correction
@@
In log messages, certain important tokens (e.g., exception information) should not be abstracted as variables. Otherwise, the log template exhibits the error of Broad template. To fix this error, we propose a variable abstracting correction strategy. Specifically, given an input log message and a template generated by the LLM, we first extract the tokens that are replaced with the wildcard “¡*¿” in the generated template. Next, we check whether these tokens contain or come after key tokens (e.g., “Exception”, “failed”, and “interrupted”) that reflect system status and are critical for engineers. If so, we use a designed correcting prompt to correct the generated template. As illustrated in Figure 5b, the LLM initially abstracted the phrase ”Could not read from stream” as a variable and replaced it with the wildcard “¡*¿”. However, because this phrase comes after the key token “IOException”, it should not be abstracted. By prompting the LLM that the phrase may not be variables, the LLM correctly identified the phrase as constants, thereby correcting the error in the originally generated template.
4. Experimental Setup
4.1. Research Questions
@@
  • RQ1: How effective is AdaParser in parsing logs?
@@
  • RQ5: How efficient is AdaParser in parsing large-scale logs?
4.2. Datasets
Our experiments are conducted using Loghub-2.0 (Jiang et al., 2023), a collection of large-scale annotated datasets for log parsing based on Loghub (Zhu et al., 2023). Loghub-2.0 includes ground-truth templates for 14 log datasets in Loghub from various types of systems. On average, each dataset within Loghub-2.0 contains 3.6 million log messages, all labeled with ground-truth log templates. Additionally, the total number of log templates across these datasets amounts to roughly 3,500.
Table 1. Accuracy comparison between AdaParser and baselines (%).
@@ 4.2. Datasets
4.3. Evaluation Metrics
Following recent studies (Khan et al., 2022; Jiang et al., 2023; JIANG et al., 2023), we used the following four metrics in our experiments:
4.3.1. Grouping Accuracy (GA)
GA is first used by Zhu et al. (Zhu et al., 2019), which assesses the ability to correctly group log messages belonging to the same template. GA is defined as the ratio of correctly grouped log messages to the total number of log messages. A log message is regarded as correctly grouped if and only if its template aligns with the same set of log messages as that of the ground truth.
4.3.2. F1 score of Grouping Accuracy (FGA)
FGA proposed by Jiang et al. (Jiang et al., 2023) is a template-level metric that focuses on the proportion of correctly grouped templates rather than log messages. Specifically, let $N_{g}$ be the actual correct number of templates in the ground truth, and $N_{p}$ be the number of templates that are generated by a log parser. If $N_{c}$ is the number of templates that are correctly parsed by the log parser, we define the Precision of Grouping Accuracy (PGA) as $\frac{N_{c}}{N_{p}}$ and the Recall of Grouping Accuracy (RGA) as $\frac{N_{c}}{N_{g}}$. Then, we calculate FGA as their harmonic mean, i.e., $\frac{2\times PGA\times RGA}{PGA+RGA}$.
4.3.3. Parsing Accuracy (PA)
PA utilized by Dai et al. (Dai et al., 2020) assesses the ability to correctly extract the static templates and dynamic variables for each log message, which is essential for downstream tasks, such as anomaly detection (Du et al., 2017). PA is defined as the ratio of correctly parsed log messages to the total number of log messages. A log message is regarded as correctly parsed if and only if all tokens of templates and variables are correctly identified.
4.3.4. F1 score of Template Accuracy (FTA)
Similar to FGA, FTA proposed by Khan et al. (Khan et al., 2022) is a template-level metric that is calculated based on the proportion of correctly identified templates. It is computed as the harmonic mean of Precision and Recall of Template Accuracy. Differently, a template is regarded as correctly identified if and only if log messages of the parsed template share the same ground-truth template and all tokens of the template are the same as those of the ground-truth template.
4.4. Baselines
In line with recent benchmark studies (Khan et al., 2022; Jiang et al., 2023), we select five open-source and state-of-the-art log parsers for comparison with our method. The first two, AEL (Jiang et al., 2008) and Drain (He et al., 2017), are chosen due to their superior performance among all syntax-based log parsers. We also choose two current top-performing semantic-based log parsers, UniParser (Liu et al., 2022) and LogPPT (Le and Zhang, 2023b). Furthermore, We incorporate the latest LLM-based log parser, LILAC (JIANG et al., 2023), which has shown the highest parsing accuracy on Loghub-2.0.
4.5. Environment and Implementation
We conducted experiments on a GPU Server equipped with an NVIDIA A100 GPU since UniParser and LogPPT require GPU resources to perform log parsing. The default LLM in AdaParser is set to ChatGPT (OpenAI, 2023) (gpt-3.5-turbo-0125), primarily due to its popularity in recent research (Le and Zhang, 2023a; Xu et al., 2024b; JIANG et al., 2023). ChatGPT was accessed via the OpenAI API, with both temperature and seed set to 0 to ensure the deterministic output for the same query, thereby enhancing reproducibility. Furthermore, we employ different LLMs to assess the generalizability of AdaParser.
To ensure fair evaluation, we adhere to the LLM-based baseline (i.e., LILAC) by default, utilizing the first 20% of the logs from each dataset as available logs. Candidates are selected using the hierarchical candidate sampling algorithm, with the default number of candidates and demonstrations set to 32 and 3, respectively. Furthermore, we vary the available logs with different ratios to evaluate the adaptability of AdaParser. For all experiments that exhibit randomness, we repeat them five times and report their mean as the final result to mitigate potential random bias.
Figure 6. Robustness comparison between AdaParser and baselines (%).
5. Evaluation Results
5.1. RQ1: How effective is AdaParser in parsing logs?
@@
5.1.1. Accuracy
Accuracy is the most critical indicator to evaluate the effectiveness of log parsers. We utilize the default settings of all methods and apply them to all log datasets. The results are shown in Table 1 with the best results for each metric on each dataset marked in bold. Metrics for AEL on the Spark dataset are denoted as “–” since it fails to complete the parsing process within a reasonable time (i.e., 12 hours), following previous works (Jiang et al., 2023; JIANG et al., 2023).

### v2

@@ III-B Tree-based Log Parser
In practical applications, the number of log templates is significantly smaller than the number of log messages [28, 17]. For instance, the Loghub-2.0 dataset [28] contains over 50 million logs but fewer than 3,500 templates. This disparity makes it inefficient to parse each log message individually using LLMs. To improve efficiency, AdaParser employs a tree-based parser, using a trie-based parsing tree to store templates generated by the LLM and quickly match templates for new log messages. Each intermediate node in the tree represents a token from log templates, with the wildcard “<*>” matching sequences of tokens of any length. The leaf node in the tree represents a complete log template, formed by concatenating all tokens along the path from the root to the leaf.
@@
When a new log message arrives, AdaParser tokenizes it and matches the tokens against the parsing tree. If the matching process reaches a leaf node, it indicates that the template stored at the leaf node exactly matches the log message, and AdaParser directly returns the template without further LLM querying. If the matching process terminates at an intermediate node, the templates within the subtrees of the intermediate node form a list of relevant templates, denoted as $[T_{1},T_{2},...,T_{n}]$, which share the same prefixes as the input log message. AdaParser then utilizes the SG-ICL-based Parser (Sec. III-C) to parse the log message and generates a corrected template $T_{c}$ through the Template Corrector (Sec. III-D). Subsequently, the parsing tree is updated based on the similarity between $T_{c}$ and the relevant templates. Specifically, we calculate the similarity between $T_{c}$ and each relevant template $T_{i}$ in $[T_{1},T_{2},...,T_{n}]$. We split them into lists of tokens, denoted as $L_{c}$ and $L_{i}$. The similarity between $L_{c}$ and $L_{i}$ is defined as follows:
$sim(T_{c},T_{i})=\frac{2\times len(LCS(L_{c},L_{i}))}{len(L_{c})+len(L_{i})}$ (1)
@@
where LCS is the longest common subsequence of the two templates. If the similarity exceeds a predefined threshold (e.g., 0.8 in our implementation), it suggests that $T_{c}$ and $T_{i}$ are highly similar and likely derive from the same ground-truth template. AdaParser will merge $T_{c}$ and $T_{i}$ by replacing the different tokens in the path of $T_{i}$ with the wildcard “<*>”. If the similarity is below the threshold, it means that $T_{c}$ has a low correlation with $T_{i}$. However, templates with low similarity can derive from the same ground-truth template. For instance, in the Linux dataset from Loghub-2.0, the similarity between the log template “apmd shutdown succeeded” and relevant templates “ntpd shutdown succeeded”, “crond shutdown succeeded”, and “xinetd shutdown succeeded” is 0.67, which is below the threshold. Nonetheless, they all share the same ground-truth template “<*> shutdown succeeded”. To address this issue, we further calculate the number of different tokens in the same position among these templates with the same similarity. If the number exceeds a predefined threshold (e.g., 5 in our implementation), AdaParser merges $T_{c}$ and $T_{i}$. Otherwise, AdaParser inserts $T_{c}$ into the parsing tree.
@@
Figure 4: The workflow of SG-ICL-based parser. After step $t$, the message with the template generated by the LLM is added to the candidate set and then used as a demonstration at step $t+1$.
III-C SG-ICL-based Parser
@@
The ICL paradigm initializes a candidate set by sampling a set of log messages and corresponding templates from the historical log data. LLMs can leverage system-specific features within demonstrations from the candidate set to facilitate more accurate log parsing. We utilize the hierarchical sampling algorithm [25] to extract a set of log messages from historical log data as the candidate set. However, as discussed in Sec. II-C, sufficient log data is not always available. Therefore, AdaParser supports initializing an empty candidate set, which is incrementally populated by SG-ICL.
@@
Figure 4 illustrates the workflow of SG-ICL-based parser. At step 00, if the candidate set is empty, we utilize zero-shot prompting to generate templates. At step $t$, when a new log message arrives, we first select similar demonstrations from the candidate set. Several studies have shown that demonstrations similar to the query can help LLMs better understand the desired behavior [39, 40, 37]. Hence, we use a kNN-based demonstration selection algorithm that is computationally efficient in practice. Specifically, for a queried log message $q$ and all candidate log messages $c_{i}$, we tokenized $q$ and $c_{i}$ into lists of tokens using delimiters such as punctuations and whitespace. Then, we calculate the LCS similarity $sim(q,c_{i})$ as shown in Eq. 1 and select the top-k most similar logs from the candidate set as demonstrations. These logs, having similar tokens and special formats to the queried log, assist the LLM in understanding the semantics of the queried log.
@@
These demonstrations are then assembled into the parsing prompt. Recent work [40] indicated that LLMs are prone to be influenced by the demonstrations that are closer to the query. Hence, we arrange the demonstrations in ascending order of similarity to the queried log. This is based on the intuition that demonstrations with higher similarity may contain more information related to the queried log. AdaParser then utilizes this parsing prompt to query the LLM and obtain the generated templates. Subsequently, a new example, i.e., the queried log message paired with the generated template, is added to the candidate set, facilitating continuous adaptation to evolving log data. At step $t+1$, the new example from the previous step $t$ becomes part of the candidate set and can be selected as a demonstration to help parse upcoming log messages.
@@
Figure 5: Two examples of self-correction. Template matching correction and variable abstracting correction are used to correct the two types of parsing errors. The corrected parts are marked in red.
III-D Template Corrector
Despite the considerable empirical success of LLMs, these models are not devoid of flaws. They occasionally exhibit undesired and inconsistent behaviors, such as generating seemingly convincing but inaccurate “hallucinations” and failing to trustfully adhere to rules and constraints [41]. In particular, when applied to log parsing, LLMs are vulnerable to two types of parsing errors, as detailed in Sec. II-C. These errors can undermine parsing accuracy and pose obstacles for downstream tasks. A prevailing strategy to correct these undesired behaviors in LLMs involves learning from feedback, analogous to the typical human learning process where individuals refine their behaviors through cycles of trial, error, and correction. When making mistakes, humans often acquire feedback either from others or through self-reflection [42, 43]. Such feedback provides valuable insights that enable humans to correct mistakes and modify their behavior accordingly.
III-D1 Template Matching Correction
@@
A correct template must accurately match the corresponding log message, otherwise, the template suffers from the error of Plausible template. To fix this error, we propose a template matching correction strategy. Specifically, given a log message and a template generated by the LLM, we first verify whether the generated template exactly matches the input log message. This verification involves transforming the generated template into a regular expression (e.g., replace “<*>” as “.*”) to check if it matches the input log message exactly. If the regular expression fails to match the input log message, we employ a designed correcting prompt to rectify the template. As illustrated in Figure 5a, the LLM initially parsed the token “0x300sent” incorrectly, treating it as a variable “0x300” and a constant “sent”. This misinterpretation led to a mismatch between the input log message and the generated log template. After querying the LLM to revise the log template, the LLM recognized that “0x300sent” should be treated as a variable and abstracted it with the wildcard “<*>”, thereby successfully correcting the error in the initially generated template.
III-D2 Variable Abstracting Correction
@@
In log messages, certain key tokens (e.g., exception information) should not be abstracted as variables, otherwise, the log template exhibits the error of Broad template. To fix this error, we propose a variable abstracting correction strategy. Specifically, given an input log message and a template generated by the LLM, we first extract the tokens that are replaced with the wildcard “<*>” in the generated template. Next, we check whether these tokens contain or come after key tokens (e.g., “Exception”, “failed”, and “interrupted”) that reflect system status and are critical for engineers. If so, we use a designed correcting prompt to correct the generated template. As illustrated in Figure 5b, the LLM initially abstracted the phrase “Could not read from stream” as a variable and replaced it with the wildcard “<*>”. However, because this phrase comes after the key token “IOException”, it should not be abstracted. By prompting the LLM that the phrase may not be variables, the LLM correctly identified the phrase as constants, thereby correcting the error in the originally generated template.
IV Experimental Setup
IV-A Research Questions
@@
  • RQ1: How effective is AdaParser?
@@
  • RQ5: How efficient is AdaParser?
IV-B Datasets
Our experiments are conducted using Loghub-2.0 [28], a collection of large-scale annotated datasets for log parsing based on Loghub [44]. Loghub-2.0 includes ground-truth templates for 14 log datasets in Loghub sourced from various types of systems, such as distributed systems, supercomputers, operating systems, and server applications. On average, each dataset within Loghub-2.0 contains 3.6 million log messages, all labeled with ground-truth log templates. Additionally, the total number of log templates across these datasets amounts to roughly 3,500.
IV-C Evaluation Metrics
Following recent studies [45, 28, 25], we used the following four metrics in our experiments:
IV-C1 Grouping Accuracy (GA)
GA [46] is a log-level metric that assesses the ability to correctly group log messages belonging to the same template. It is defined as the ratio of correctly grouped log messages over all log messages. A log message is regarded as correctly grouped if and only if its template has the same group of log messages as the oracle.
IV-C2 F1 score of Grouping Accuracy (FGA)
FGA [28] is a template-level metric that measures the proportion of correctly grouped templates. It is computed as the harmonic mean of precision and recall of grouping accuracy, where the template is considered as correct if log messages of the predicted template have the same group of log messages as the oracle.
IV-C3 Parsing Accuracy (PA)
PA is a log-level metrics [12] that measures the correctness of extracted templates and variables. It is defined as the ratio of correctly parsed log messages to the total number of log messages. A log message is regarded as correctly parsed if and only if all tokens of templates and variables are identical with the oracle.
IV-C4 F1 score of Template Accuracy (FTA)
FTA is a template-level metric that measures the proportion of correctly identified templates. It is computed as the harmonic mean of precision and recall of Template Accuracy. A template is regarded as correctly identified if and only if log messages of the parsed template share the same ground-truth template and all tokens of the template are the same as those of the ground-truth template.
IV-D Baselines
In line with recent benchmark studies [45, 28], we select five open-source and state-of-the-art log parsers for comparison with our method. The first two, AEL [47] and Drain [11], are chosen due to their superior performance among all syntax-based log parsers. We also choose two current top-performing semantic-based log parsers, UniParser [13] and LogPPT [14]. Furthermore, We incorporate the latest LLM-based log parser, LILAC [25], which has shown the highest parsing accuracy on Loghub-2.0. To ensure a fair comparison, we utilize the implementations of all baselines from their replication repositories, choosing the default settings or hyperparameters.
IV-E Environment and Implementation
We conducted experiments on a GPU Server equipped with an NVIDIA A100 GPU since UniParser and LogPPT require GPU resources to perform log parsing. The default LLM in AdaParser is set to ChatGPT [48] (gpt-3.5-turbo-0125), primarily due to its popularity in recent research [24, 26, 25]. ChatGPT was accessed via the OpenAI API, with both temperature and seed set to 0 to ensure the deterministic output for the same query, thereby enhancing reproducibility. Furthermore, we employ different LLMs to assess the generalizability of AdaParser in RQ4.
To ensure fair evaluation, we adhere to the LLM-based baseline (i.e., LILAC) by default, utilizing the first 20% of the logs from each dataset as available logs. Candidates are selected using the hierarchical candidate sampling algorithm, with the default number of candidates and demonstrations set to 32 and 3, respectively. Furthermore, we vary the available logs with different ratios to evaluate the adaptability of AdaParser in RQ2. We have implemented AdaParser in Python and integrated it into previous benchmarks [46, 45, 28], allowing for a fair comparison between AdaParser and all baselines in the same framework. For all experiments that exhibit randomness, we repeat them five times and report their mean as the final result to mitigate potential random bias.
TABLE I: Accuracy comparison between AdaParser and baselines (%).
@@ IV-E Environment and Implementation
Figure 6: Robustness comparison between AdaParser and baselines (%).
V Evaluation Results
V-A RQ1: How effective is AdaParser?
@@
V-A1 Accuracy
Accuracy is the most critical indicator to evaluate the effectiveness of log parsers. We utilize the default settings of all methods and apply them to all log datasets. The results are shown in Table I with the best results for each metric on each dataset marked in bold. Metrics for AEL on the Spark dataset are denoted as “–” since it fails to complete the parsing process within a reasonable time (i.e., 12 hours), following previous works [28, 25].